OpenStudy (anonymous):

Find the limit as n approaches infinity of (1+ 2/n) ^ (1/n)

OpenStudy (anonymous):

I know it is 1 but I am not sure how to show it...?

OpenStudy (anonymous):

make a guess i bet you are right

OpenStudy (anonymous):

Does it equal one just because any number to the power of 0 to 1?

OpenStudy (anonymous):

is not to*

OpenStudy (anonymous):

Is there any way I can show my work?

OpenStudy (anonymous):

oh no, it is definitely not one

OpenStudy (anonymous):

Howw?

OpenStudy (anonymous):

ok you win it is one i read it wrong q

OpenStudy (anonymous):

Nah it's good, how can I show this work though?

OpenStudy (anonymous):

take the log to get the variable out of the exponent and start with \[\frac{1}{n}\ln(1+\frac{2}{n})\]

OpenStudy (anonymous):

Lhospitals?

OpenStudy (anonymous):

take the limit of that one, you will get \(0\) so your answer is \(e^0=1\)

OpenStudy (anonymous):

yeah you can do it using l'hopital

OpenStudy (anonymous):

how did you get the e^0 part?

OpenStudy (anonymous):

oh I got it. thanks

OpenStudy (anonymous):

because the first step was taking the log

OpenStudy (anonymous):

yw